# gpu_ext: Extensible OS Policies for GPUs via eBPF

### v1

@@ Abstract
Performance in modern GPU-centric systems depends increasingly on resource management policies, such as memory placement, scheduling, and observability. However, a one-size-fits-all policy performs poorly across diverse workloads. Existing approaches present a tradeoff: user-space runtimes offer programmability but lack cross-tenant visibility and fine-grained hardware control, while OS kernel modification introduce complexity and safety risks. To address this, we argue that the GPU driver and device layer must serve as an extensible OS policy interface. The emerging eBPF offers a possibility, but naively transplanting host-side eBPF is insufficient: it cannot observe critical device-side events, and directly injecting policy code into GPU kernels affects safety and efficiency.
We present gpu_ext, an eBPF-based policy runtime that treats the GPU driver and device as a programmable OS subsystem. gpu_ext extends GPU drivers to expose safe hooks and introduces a device-side eBPF runtime that executes verified policy logic within GPU kernels, enabling coherent, application-transparent policies. Evaluation on realistic workloads, including inference, training, and vector search, shows that gpu_ext improves throughput by up to and reduces tail latency by up to with low overhead, without modifying applications or restarting drivers.
@@ 1 Introduction
Unfortunately, existing GPU resource management policy solutions force developers into tradeoffs between user-space methods that lack global visibility and control over system-level GPU behaviors, and kernel-level implementations that introduce complexity and safety risks. Concretely, user-space frameworks [ng2023paella, shen2025xsched, gim2025pie, chen2025ktransformers] offer programmability, but lack cross-application visibility and management of low-level GPU mechanisms (e.g. thread timeslices and replayable page faults [fan2025gpreempt, nvidia-open-gpu]). These approaches bind to specific runtime frameworks, requiring application-level code changes and limiting portability and reuse. Conversely, kernel-based resource management approaches [kato2011timegraph, kato2012gdev, kernelintercept, fan2025gpreempt, wang2024gcaps] provide visibility and control of GPU hardware states, their direct modification on static, vendor-specific kernel modules making them difficult to deploy and maintain, and posing stability concerns. GPU profiling frameworks [cupti, villa2019nvbit, neutrino], including our prior workshop version [egpu], enable kernel-level observation but lack runtime programmability and driver integration, limiting adaptive resource management.
To address this, we argue that GPU resource management requires an extensible OS policy interface for GPU providing safe, flexible and transparent programmability. The GPU driver layer is positioned to offer global visibility, enable cross-tenant coordination, and provide fine-grained management of privileged low-level GPU mechanisms without requiring application modifications. Recent CPU-side frameworks, such as eBPF, have demonstrated the feasibility of dynamic, kernel-level policy programmability with safety guarantees [bachl2021flow, Zhong22, jia2023programmable, schedext-docs, zussman2025cache_ext]. Inspired by its success, we propose treating GPU resource management as a first-class programmable OS subsystem. However, naively extending host-only eBPF remains insufficient because scheduling and memory management decisions such as thread-block dispatch, warp-level execution scheduling, and memory synchronization occur within GPU hardware and remain invisible and uncontrollable from the host. Thus, a practical solution must expose programmable policy hooks within GPU drivers and execute verified policy logic inside GPU kernels.

### v2

@@ Abstract
Performance in modern GPU-centric systems increasingly depends on resource management policies, including memory placement, scheduling, and observability. However, uniform policies typically yield suboptimal performance across diverse workloads. Existing approaches present a tradeoff: user-space runtimes provide programmability and flexibility but lack cross-tenant visibility and fine-grained control of hardware resources; meanwhile, modifications to the OS kernel introduce significant complexity and safety risks. To address this, we argue that the GPU driver and device layer should provide an extensible OS interface for policy enforcement. While the emerging eBPF technology shows potential, directly applying existing host-side eBPF is insufficient because they lack visibility and control into critical device-side events, and directly embedding policy code into GPU kernels could compromise safety and efficiency.
We propose gpu_ext, an eBPF-based runtime that treats the GPU driver and device as a programmable OS subsystem. gpu_ext extends GPU drivers by exposing safe programmable hooks and introduces a device-side eBPF runtime capable of executing verified policy logic within GPU kernels, enabling coherent and transparent policies. Evaluation across realistic workloads including inference, training, and vector search demonstrates that gpu_ext improves throughput by up to and reduces tail latency by up to , incurring minimal overhead, without modifying or restarting applications.
@@ 1 Introduction
Unfortunately, existing programmable GPU resource management policy solutions force developers into tradeoffs between user-space methods that lack global visibility and control over system-level GPU behaviors, and kernel-level implementations that introduce complexity and safety risks. Concretely, user-space frameworks [ng2023paella, shen2025xsched, gim2025pie, chen2025ktransformers] offer programmability, but lack cross-application visibility and management of low-level GPU mechanisms (e.g. thread timeslices and replayable page faults [fan2025gpreempt, nvidia-open-gpu]). These approaches bind to specific runtime frameworks, requiring application-level code changes and limiting portability and reuse. Conversely, kernel-based resource management approaches [kato2011timegraph, kato2012gdev, kernelintercept, fan2025gpreempt, wang2024gcaps] provide visibility and control of GPU hardware states, their direct modification on static, vendor-specific kernel modules making them difficult to deploy and maintain, and posing stability concerns. GPU profiling frameworks [cupti, villa2019nvbit, neutrino], including our prior workshop version [egpu], enable kernel-level observation but lack runtime programmability and driver integration, limiting adaptive resource management.
To address this, we argue that GPU resource management requires an OS policy interface for GPU providing safe, flexible and transparent programmability. The GPU driver layer is positioned to offer global visibility, enable cross-tenant coordination, and provide fine-grained management of privileged low-level GPU mechanisms without requiring application modifications. Recent CPU-side frameworks, such as eBPF, have demonstrated the feasibility of dynamic, kernel-level policy programmability with safety guarantees [bachl2021flow, Zhong22, jia2023programmable, schedext-docs, zussman2025cache_ext]. Inspired by its success, we propose treating GPU resource management as a first-class programmable OS subsystem. However, naively extending host-only eBPF remains insufficient because scheduling and memory management decisions such as thread-block dispatch, warp-level execution scheduling, and memory synchronization occur within GPU hardware and remain invisible and uncontrollable from the host. Thus, a practical solution must expose programmable policy hooks within GPU drivers and execute verified policy logic inside GPU kernels.